OpenStudy (anonymous):

Evaluate the expression below when x = -1. 2x2 − 5 x + 9

OpenStudy (anonymous):

somebody help me please

OpenStudy (anonymous):

please help me and sheppard18 these questions is what we are depending on to graduate

OpenStudy (anonymous):

??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

hello??

Nnesha (nnesha):

replace x by -1

OpenStudy (anonymous):

is that the answer

Nnesha (nnesha):

nope

OpenStudy (anonymous):

whats the answer

OpenStudy (confluxepic):

Just substitute it.

Nnesha (nnesha):

\[2(-1)^2 -5\] substitute value and then solve !

OpenStudy (anonymous):

i dont know how to substitute because we have never ever had a half decent math teacher

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[2(-1)^2 -5\] substitute value and then solve ! \(\color{blue}{\text{End of Quote}}\) like this just replace x and plug -1 now solve what is square of (-1) ?

Nnesha (nnesha):

(-1)^2 is same as -1 times -1 = what ?

OpenStudy (anonymous):

i dont know what the answer is

OpenStudy (confluxepic):

For example x=12. 3(x)+2=y 3(12)+2=y <--------------------------------- 36+2=y | 38=y | | See how I substituted x as 12 to solve the example equation? -------

Nnesha (nnesha):

what is -1 times -1 = ??

OpenStudy (confluxepic):

Arrow line messed up.

OpenStudy (anonymous):

-1?

Nnesha (nnesha):

negative times negative = positive negative times positive = negative so negative 1 times negative 1 = ?

Nnesha (nnesha):

@sheppard18 what do you think ? -1 times -1 = ??

OpenStudy (anonymous):

1?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

1 is not a choice

Nnesha (nnesha):

\[2(-1)^2 -5\] \[2(1)-5 =what ??\]

Nnesha (nnesha):

i didn't say one is the answer

OpenStudy (anonymous):

-7/8, -3/8, -1/8, -1 theses are the choices we have

OpenStudy (anonymous):

-3?

OpenStudy (confluxepic):

Example: x+2=y You know that x is 1 but you need to find out y. So you put \(\color{red}{x~as~1}\) in the equation. \(\color{red}{(1)}\)+2=y. 1+2=3 so y=3. Get it?

Nnesha (nnesha):

yes that's right! remember PEMDAs rule Parentheses exponent MD) MULTIPLY OR DIVIDE AS ) ADD OR SUBTRACT

OpenStudy (anonymous):

please just tell me the answer

OpenStudy (anonymous):

were from tug valley we dont know much

Nnesha (nnesha):

you have to learn in order to know something

OpenStudy (anonymous):

are teacher does nothing but sleep in class

OpenStudy (anonymous):

and calls the office on our class because our class is so mean

Nnesha (nnesha):

alright thanks for joining os here we will teach u no direct answers!!

OpenStudy (anonymous):

well how do i do it

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @estepp33 -3? \(\color{blue}{\text{End of Quote}}\) yes right

OpenStudy (anonymous):

what about -3

Nnesha (nnesha):

substitute given x value which is -1 \[\huge\rm 2(-1)^2 -5 \] take square of -1 which is 1 (-1 times -1 = 1) \[2(1)-5 = ???\]

OpenStudy (anonymous):

that comes out to be -3 what do we do with the -3

Nnesha (nnesha):

that's the answer of first one

Nnesha (nnesha):

i guess your question is \[\frac{ 2x^2 -5 }{ x+9 }\] right it's a fraction ?

OpenStudy (anonymous):

it will be a fraction or a whole numbner

OpenStudy (anonymous):

yes its a fraction i need help with the bottom now

Nnesha (nnesha):

so why you didn't tell me ?? use */* for fraction so we know it's a fraction

Nnesha (nnesha):

i thought those are 2 different expression

Nnesha (nnesha):

\[\huge\rm \frac{ 2\color{blue}{x^2} -5 }{ \color{blue}{x}+9 }\] just like we did before replace x by -1

OpenStudy (anonymous):

would the answer be -3/8

Nnesha (nnesha):

that's right

OpenStudy (anonymous):

i got 8

Nnesha (nnesha):

yes the denominator is 8

OpenStudy (anonymous):

can u help me with this question The simplest form of the expression 4m−17m2−16+3m−11m2−16 has in the numerator and in the denominator.

Nnesha (nnesha):

and how do i know what is in the numerator and what is in the denominator use the equation and type ur question

OpenStudy (anonymous):

i dont know

Nnesha (nnesha):

what ?

Nnesha (nnesha):

\[\huge\rm 4m−17m^2−16+3m−11m^2−16\] like this ?

OpenStudy (anonymous):

4m-17/m^2-16+3m-11/m^2-16

OpenStudy (anonymous):

that is the fraction ^

OpenStudy (anonymous):

4m-17/m^2 this is the fraction

Nnesha (nnesha):

\[\huge\rm \frac{ 4m-17 }{ m^2 } -\frac{ 16+3m-11 }{ m^2 -16 }\] like this ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

m^2-16 on both bottom sides

Nnesha (nnesha):

\[\huge\rm \frac{ 4m-17 }{ m^2 -16} -\frac{ 16+3m-11 }{ m^2 -16 }\] looks good ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no

Nnesha (nnesha):

alright type your question please

OpenStudy (anonymous):

4m-17/m^2-16 + 3m-11/m^2-16

OpenStudy (anonymous):

the minus sign should be plus and the 16 on top on the right shouldnt be there

Nnesha (nnesha):

\[\huge\rm \frac{ 4m-17 }{ m^2 -16} +\frac{ 3m-11 }{ m^2 -16 }\] find common denominator? what is common denominator ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

??

Nnesha (nnesha):

1not just 16 m^2 -16 should be together \[\huge\rm \frac{ }{m^2 -16 }\] so m^2 and -16= common denominator

OpenStudy (anonymous):

ok what do i do

Nnesha (nnesha):

now bec denominator of both fraction are same as common denominator so just carry down the value of numerator \[\huge\rm \frac{ 4m-17+3m-11 }{ m^2 -16 }\] now combine like terms 4m-17+3m -11= what ?

OpenStudy (anonymous):

-21

Nnesha (nnesha):

how did yo uget -21 ?

OpenStudy (anonymous):

i done 4-17+3-11

Nnesha (nnesha):

what are LIKe terms ?

Nnesha (nnesha):

you can't combine 4m with -17 bec there are not the same

OpenStudy (anonymous):

3 and 4

Nnesha (nnesha):

add coefficient of same variable yes right 3m and 4m combine them

Nnesha (nnesha):

3m+4m =what ? -17-11= ?

OpenStudy (anonymous):

7 and -28

OpenStudy (anonymous):

??

Nnesha (nnesha):

\[ \huge\rm \frac{ 7m -28 }{ m^2 -16 }\] now factor 7m-28 what is common factor ?

Nnesha (nnesha):

nd apply difference of square for the denominator part \[\rm a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

what do i do with that

Nnesha (nnesha):

alright first factor the numerator 7m-28 what is common factor ?

OpenStudy (anonymous):

is it m+4

OpenStudy (anonymous):

??

Nnesha (nnesha):

for which one ?

OpenStudy (anonymous):

the right one

Nnesha (nnesha):

alright 7m-28 what is common in these both terms ??

OpenStudy (anonymous):

i got to go im going home

OpenStudy (anonymous):

thanks for the help